40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 30 13:43:43 2008, rob, Update, Locking, last night 
    Reply  Wed Oct 1 00:24:54 2008, rana, Update, Locking, last night 
Message ID: 1011     Entry time: Wed Oct 1 00:24:54 2008     In reply to: 1009
Author: rana 
Type: Update 
Category: Locking 
Subject: last night 
I had mistakenly left the MC boost off during my FAST investigations. The script is now restored.

The ISS is still saturating with gains higher than -5 dB. We need to request a PeterK / Stefan consult in the morning.

Also found the MZ gain down at -10 dB around midnight - need an alarm on that value.
ELOG V3.1.3-